OpenStudy (anonymous):

Reduce to lowest terms: 1. x-2/3x-6 Write as a single fraction in lowest terms: 1. x/4+x/5 2. x+6/4 multiplied by 2/3 3. 1/8 divided by 8/x

jimthompson5910 (jim_thompson5910):

To reduce \(\Large \frac{x-2}{3x-6}\), factor out 3 from the expression 3x-6 to get 3(x-2). I'm sure at this point you see what cancels

OpenStudy (bohotness):

Do you known how to reduce to the lower term?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Can u help me witht he other problems as well, Jim?

jimthompson5910 (jim_thompson5910):

what did you get for your answer?

OpenStudy (anonymous):

x-2

jimthompson5910 (jim_thompson5910):

\[\Large \frac{x-2}{3x-6} = \frac{x-2}{3(x-2)}\] what terms cancel?

OpenStudy (anonymous):

x-2 and 3x

jimthompson5910 (jim_thompson5910):

the "x-2" terms are the only things that cancel

jimthompson5910 (jim_thompson5910):

\[\Large \frac{x-2}{3x-6} = \frac{x-2}{3(x-2)}\] \[\Large \frac{x-2}{3x-6} = \frac{\cancel{x-2}}{3\cancel{(x-2)}}\] \[\Large \frac{x-2}{3x-6} = \frac{1}{3}\]

OpenStudy (anonymous):

ohhh I get it now